# Uncertainties from metrology in the integrated luminosity measurement with the updated design of the very forward region of a detector at CEPC

### v1

@@ Abstract
At an $e^{+}e^{-}$ collider machine-detector interface includes, among others, a calorimeter dedicated for precision measurement of the integrated luminosity at a per mill level or better. Usually, such a device is placed at the outgoing beams, to keep the spatial symmetries of the head-on collisions at accelerators with a non-zero crossing angle. At CEPC it is currently proposed that the luminometer is placed on the z-axis. We review a feasibility of precision measurement of the integrated luminosity at the $Z^{0}$ pole, from the point of view of systematic effects arising from metrology uncertainties, either concerning the luminometer position and alignment or the beam properties, with the luminometer centered around the z-axis.
@@ 1 Introduction
To meet precision requirements of the CEPC physics program, relative uncertainty of the integrated luminosity measurement should be of order of $10^{-4}$ at the $Z^{0}$ pole CEPC_CDR . The method of integrated luminosity measurement at CEPC, as well as the machine parameters, detector concept, machine-detector interface (MDI) and physics performance, is described in CEPC_CDR . A distinctive feature of the current MDI design at CEPC is that the luminometer is proposed to be placed at the z-axis. Although the impact of metrology on the integrated luminosity measurement at CEPC has been studied in JINST with the luminometer conventionally placed at the outgoing beams (s-axis), it is worth revisiting for the newly proposed geometry, since the integrated luminosity precision critically depends on the luminometer’s position, as well as on the way of counting.
@@ 2 Forward region at CEPC
The MDI region of CEPC covers the area of 6 m from the interaction point (IP) in both directions, along the z-axis. The accelerator components inside the detector without shielding are within a conical space with an opening angle of 118 mrad. The crossing angle in the horizontal plane between the colliding beams at the IP is 33 mrad. The final focus length of colliding beams is 2.2 m CEPC_CDR . It is proposed that the luminometer at CEPC should be placed at 950 mm distance from the interaction point, covering the polar angle region between 30 mrad and 105 mrad, which corresponds to the luminometer aperture of 28.5 mm for the inner radius and 100 mm for the outer. The fiducial volume of luminometer should be between 53 mrad and 79 mrad. The luminometer will be placed half way to the tracking volume longitudinally. The layout of the MDI region at CEPC is illustrated in Fig. 1 JINST . Shower leakage from the outer edge of the luminometer has been studied and proven to be negligible after absorption by a 5 mm iron filter positioned around the luminometer CEPC_CDR . However, this has been done under assumption that the luminometer will be placed around the outgoing beam and it is yet to be proven for the luminometer placed around the z-axis.
Figure 1: Layout of the MDI region at CEPC.
@@ 3 Integrated luminosity measurement and systematic uncertainties
Integrated luminosity measurement is a counting experiment conventionally based on the low angle Bhabha scattering (LAB). It is defined as:
@@
where $N_{Bh}$ is Bhabha count in the certain phase space of parameters and $\sigma_{Bh}$ the SAB cross-section.
To control integrated luminosity at the required level of $10^{-4}$ at the $Z^{0}$ pole, both Bhabha count and theoretical cross-section should be known with the same precision. The Bhabha cross-section scales with the center-of-mass energy as FCAL_JINST :
@@
where $\alpha_{em}$ is the QED constant and $\theta$ the polar angle of the emitted Bhabha particles. Uncertainty of $\mathrm{10^{-4}}$ of the cross-section implies that the available center-of-mass energy should be known with $\Delta(\sqrt{s})\leq$ 5 MeV. There is an ongoing work at future Higgs factories nesto to check the feasibility of $\sqrt{s}$ determination using di-muon production reconstructed in the central tracker.
As discussed in Stahl , many s-channel processes far from the resonance will have the same cross-section dependence on the center-of-mass energy, leading to effective cancellation of the uncertainties in the cross-section measurement. It is up to individual cross-section measurement analysis to determine if uncertainty of the integrated luminosity measurement caused by the uncertainty of $\sqrt{s}$ is an issue.
Placement of the luminometer on the z-axis with the CEPC crossing angle of 33 mrad leads to the reduction in Bhabha count for $\sim 70$% in comparison to the count with the luminometer placed on the s-axis. Since the overall Bhabha count in 16 $\mathrm{ab}^{-1}$ of integrated luminosity expected at the $Z^{0}$ pole will be of order of $10^{12}$ for the detector placed on the s-axis, its displacement to the z-axis with the same luminometer apertures will contribute to the relative statistical uncertainty of the count as $\mathrm{1.2\cdot 10^{-6}}$. Hitmap in the first plane of the luminometer will look different, as can be seen on Fig. 2, with the expected loss of the axial symmetry.
Further we discuss feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole, considering that each systematic effect individual contribution to the uncertainty of the integrated luminosity is of order of $\mathrm{10^{-4}}$. We have assumed event selection in polar angle acceptance to be asymmetric between the left and right arms of the detector. In JINST , at the one side of the luminometer, the full fiducial volume is considered, while at the other side the inner radial acceptance is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm, as it has been done at OPAL. This has been done randomly to the left (L) and right (R) sides of the luminometer, on event by event basis. For symmetric event selection we consider the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way. Fig. 3 provides an illustrative depiction of the s- and z-axes configuration in collisions with the crossing angle.
@@ 3.1 Uncertainties from mechanics and positioning
Systematic uncertainties from detector and beam related uncertainties have been quantified through a simulation study, assuming $2\cdot 10^{7}$ Bhabha scattering events generated using BHLUMI V4.04 Bhabha event generator BHLUMI , at the $Z^{0}$ production threshold. Final state particles are generated in the polar angle range from 30 mrad to 100 mrad that is within a couple of mrad margin outside of the detector fiducial volume, to allow events with non-collinear final state radiation to contribute. Initial state radiation is also included. The effective Bhabha cross-section in this angular range is $\sim$50 nb.
Considered detector-related uncertainties arising from manufacturing, positioning and alignment of the luminometer are:
@@
    Maximal uncertainty of the luminometer inner radius, $\Delta r_{in}$,
@@
    RMS of the Gaussian spread of the measured radial shower position with respect to the true hit position, $\sigma_{r}$ (i. e. by placing a tracker plane in front of the luminometer),
@@
    Maximal absolute uncertainty of the longitudinal distance between left and right halves of the luminometer, $\Delta l$, where both halves are shifted for ±$\pm\frac{\Delta l}{2}$ with respect to the interaction point,
@@
    RMS of the Gaussian distribution of mechanical fluctuations of the luminometer position with respect to the interaction point (IP), caused by vibrations and thermal stress, radial, $\sigma_{x_{IP}}$ and axial, $\sigma_{z_{IP}}$.
Considered deviations are maximal, as we assumed $10^{-4}$ contribution to the relative uncertainty of integrated luminosity from each individual effect. Figures 4-8 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis. Limits of mechanical parameters are listed in table 1.
@@
Figure 4: (a) Variations of the inner aperture, $\Delta r_{in}$, for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), for detector placed on the s- and z-axis; (b) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed ot the z-axis.
Figure 5: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS, of the Gaussian spread of the measured radial shower position with respect to the true impact position in the luminometer front plane, $\sigma_{r}$. (b) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed ot the z-axis.
Figure 6: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal absolute uncertainty, $\Delta l$, of the longitudinal distance between left and right halves of the luminometer; (b) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 7: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian distribution of mechanical fluctuations of the luminometer position with respect to the IP, caused by radial vibrations and thermal stress, $\sigma_{x_{IP}}$; (b) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 8: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian distribution of mechanical fluctuations of the luminometer position with respect to the IP, caused by axial vibrations and thermal stress, $\sigma_{z_{IP}}$; (b) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@ 3.2 Beam related uncertainties
Several systematic uncertainties may arise from the uncertainties of beam properties and its delivery to the interaction point. The list of CEPC beam parameters addressed in this study is given in Table 1 Gao .
Table 1: CEPC TDR parameters at the $Z^{0}$ pole.
@@
Here we consider:
@@
    maximal permanent bias ($\Delta E$) of a single beam energy with respect to the other beam, resulting in longitudinal boost of the incoming $e^{+}e^{-}$ system with respect to the laboratory frame, and consequently, boost of the Bhabha center-of-mass system,
@@
    maximal half-width of the Gaussian of the beam energy spread ($\sigma_{E_{BS}}$), responsible for longitudinal boost on event-by-event basis, leading to the overall count loss of order of $\mathrm{10^{-4}}$,
@@
    maximal radial ($\Delta x_{IP}^{BS}$) and axial ($\Delta z_{IP}^{SY}$) IP position displacements with respect to the luminometer, caused by the finite beam sizes (former) and beam synchronization (latter),
@@
    maximal time shift in beam synchronization ($\Delta\tau_{SY}$), calculated from $\Delta z_{IP}^{SY}$, leading to the IP longitudinal displacement $\Delta z_{IP}^{SY}$.
In the same manner as in the previous chapter, results for effects 1-3 are illustrated on Figures 9-12 and the upper limits are summarized in table 2, each contributing to the relative uncertainty of the integrated luminosity at the $Z^{0}$ pole as $10^{-4}$.
Table 2: Required minimal absolute precision of mechanical and beam parameters contributing to the relative uncertainty of the integrated luminosity of $10^{-4}$ at the $Z^{0}$ pole. The average net center-of-mass energy uncertainty $\Delta E_{CM}$ limit is derived by error propagation from the Bhabha cross-section dependence on the center-of-mass energy.
@@
As can be seen from Table 2, mechanical precision of the luminometer inner aperture is the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$.
@@
Figure 10: (b) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal half-width of the BES, corresponding to a relative statistical uncertainty of $\mathcal{L}$, $\sigma_{E_{BS}}$; (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S-triangles, AS-diamonds), with the luminometer on the z-axis.
@@
The longitudinal boost of the Bhabha center-of-mass frame with respect to the laboratory frame ($\beta_{z}$) may be caused by any difference in beam energies, $\delta E$, in the form of bias $\Delta E$ of one beam with respect to the other, or as a random asymmetry caused by the beam energy spread ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), or by the initial state radiation of beams (not considered here). Limits of $\Delta E$ and $\sigma_{E_{BS}}$ given in Table 2 are derived from Figures 9(b) and 10(b) and rounded with respect to statistical size of the sample. From Table 2 it can be seen that in terms of the beam-related parameters precision, displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision, while some dependencies will be lost due to loss of spacial symmetries of Bhabha events.
Due to hard precision requirement on luminometer inner aperture, as well as to numerous and sometimes complex systematic effects in integrated luminosity measurement, like the effects caused by beam-beam interactions Voutsinas , it is worth to explore, in addition, a possibility to measure integrated luminosity with additional central processes, such as di-muon or di-photon production ECFA .

### v2

# Uncertainties from metrology in the integrated luminosity measurement with the updated design of a detector at CEPC

@@ Abstract
In order to measure integrated luminosity with a required precision of $10^{-4}$ at the $Z^{0}$ pole, proposed CEPC $e^{+}e^{-}$ collider requires a luminometer, a specially designed calorimeter placed in the very forward region to identify Bhabha scattering at low polar angles. Usually, such a device is placed at the outgoing beams, to keep the spatial symmetries of the head-on collisions at accelerators with a non-zero crossing angle. At CEPC it is currently proposed to place the luminometer on the z-axis. We review a feasibility of a measurement of the integrated luminosity at the $Z^{0}$ pole with the required precision, concerning the luminometer centered around the z-axis and the post-CDR beam properties.
@@ 1 Introduction
In order to achieve precision goals of the electroweak physics program at CEPC at the $Z^{0}$ pole, integrated luminosity $\mathcal{L}$ should be known with the total relative uncertainty of order of $10^{-4}$ CEPC_CDR . In the CEPC Conceptual Design Report CEPC_CDR , beam properties and detector concept are discussed, with particular emphasis on the machine-detector interface (MDI) and integrated luminosity measurement. The current design of MDI at CEPC has a distinctive feature with respect to the proposal in CEPC_CDR - the luminometer is proposed to be placed at the z-axis. Although the impact of metrology on the integrated luminosity measurement at CEPC with the luminometer conventionally placed at the outgoing beams (s-axis) has been discussed in JINST , it requires revisiting for the newly proposed geometry, since the integrated luminosity precision critically depends on the luminometer’s position, as well as on the event selection.
@@ 2 Forward region at CEPC
Figure 1: Schematic view of the MDI region at CEPC.
The MDI region of CEPC is proposed to longitudinally cover the area of 12 m with the centrally positioned interaction point (IP). The components of the accelerator (without shielding) will be placed in an 118 mrad conus inside the detector, as illustrated in Fig. 1 JINST . The crossing angle in the horizontal plane between the colliding beams at the IP is 33 mrad. The final focus length of colliding beams is 220 cm CEPC_CDR . It is proposed that the luminometer at CEPC should be placed at 95 cm distance from the interaction point, covering the polar angle region from 30 mrad to 105 mrad, which corresponds to the luminometer aperture of 2.85 cm for the inner and 10.0 cm for the outer radius. The fiducial volume of luminometer should be between 53 mrad and 79 mrad. To prevent the leakage of the electromagnetic showers outside of the luminometer’s edges, an iron shielding of 5 mm thickness can be placed around the luminometer. However, this has been done under assumption that the luminometer will be placed around the outgoing beams and it is yet to be proven for the luminometer placed around the z-axis, since the symmetry of signal hits with respect to the detector axis will be lost, as illustrated in Fig. 2, where the hitmap in the first plane of the luminometer is presented.
@@ 3 Integrated luminosity measurement and systematic uncertainties
Integrated luminosity measurement is conventionally based on the low angle Bhabha scattering (LABS). It is defined as:
@@
where $N_{Bh}$ is Bhabha count in the certain phase space of parameters and $\sigma_{Bh}$ is the LABS cross-section. Both Bhabha count and theoretical cross-section at the $Z^{0}$ pole should be known at the level of $10^{-4}$, in order to know the integrated luminosity with the same precision. The Bhabha cross-section scales with the center-of-mass energy as FCAL_JINST :
@@
where $\alpha_{em}$ is the QED constant and $\theta$ the polar angle of emitted Bhabha particles. Uncertainty of $\mathrm{10^{-4}}$ of the cross-section implies that the available center-of-mass energy should be known with the absolute uncertainty of $\Delta(\sqrt{s})\leq$ 5 MeV. There is an ongoing work at future Higgs factories nesto to check the feasibility of $\sqrt{s}$ determination using di-muon production reconstructed in the central tracker. However, the required precision seems to be beyond the experimental reach, but, as discussed in Stahl , many s-channel processes far from the resonance will have the same dependence of the cross-section on the center-of-mass energy, leading to effective cancellation of the uncertainties in the cross-section measurements. It is up to each individual cross-section measurement analysis to determine if uncertainty of the integrated luminosity measurement caused by the uncertainty of $\sqrt{s}$ is an issue.
Placement of the luminometer on the z-axis with the CEPC crossing angle of 33 mrad leads to the reduction in Bhabha count for $\sim 70$% in comparison to the count with the luminometer placed on the s-axis. Fig. 3 provides an illustrative depiction of the s- and z-axes configurations in collisions with the crossing angle. Since the overall Bhabha count in 16 $\mathrm{ab}^{-1}$ of integrated luminosity, which is expected at CEPC at the $Z^{0}$ pole, will be of order of $10^{12}$ for the detector placed on the s-axis, its displacement to the z-axis with the same luminometer apertures will contribute to the relative statistical uncertainty of the count as $\mathrm{1.2\cdot 10^{-6}}$.
@@
Further, feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole will be discussed, considering the uncertainty of each individual systematic effect contributes to the integrated luminosity at a level of $\mathrm{10^{-4}}$. In JINST , at one side of the luminometer the Bhabhas are counted in the full fiducial volume of the luminometer, while at the other side the inner radial acceptance of the fiducial volume is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm, as it has been done at OPAL OPAL , randomly to the left (L) and right (R) sides of the luminometer, on event by event basis. Described way of counting is called asymmetrical throughout the paper. For symmetric event selection the Bhabhas are counted in the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way.
@@ 3.1 Uncertainties from mechanics and positioning
Systematic uncertainties arising from the detector mechanics and positioning, as well as the uncertainties related to the properties of beams, have been quantified through a simulation study. We have generated, including the effects of initial (ISR) and final state radiation (FSR), $2\cdot 10^{7}$ Bhabha events scattered at low angles at the $Z^{0}$ pole, using BHLUMI V4.04 Bhabha event generator BHLUMI . The polar angle range in which the final state particles are generated is from 30 mrad to 100 mrad, to allow events modified by FSR to be detected in the luminometer. In this angular range, the effective Bhabha cross-section is $\sim$50 nb.
Considered detector-related uncertainties originating from manufacturing, positioning and alignment of the luminometer are, as in JINST :
@@
    Maximal uncertainty of the luminometer inner aperture, $\Delta r_{in}$,
@@
    RMS of the Gaussian dissipation of the radial shower position measured in the luminometer, with respect to the exact position of the Bhabha hit, $\sigma_{r}$ (measured i. e. by placing a tracker plane in front of the luminometer),
@@
    Maximal absolute uncertainty of the distance between left and right sides of the luminometer along the z-axis, $\Delta l$, where both halves are shifted equidistantly for ±$\pm\frac{\Delta l}{2}$ with respect to the interaction point,
@@
    RMS of the Gaussian distribution of luminometer fluctuations with respect to the IP, caused by vibrations and thermal stress, in radial, $\sigma_{x_{IP}}$, and axial direction, $\sigma_{z_{IP}}$.
Figures 4-8 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis.
@@
Figure 4: (a) Variations of the inner aperture, $\Delta r_{in}$, for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), for detector placed on the s- and z-axis; (b) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 5: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian spread of measured radial shower position with respect to the true one, $\sigma_{r}$. (b) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 6: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal absolute uncertainty, $\Delta l$, of the longitudinal distance between left and right luminometer sides; (b) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 7: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of a Gaussian distribution of radial fluctuations of the luminometer position with respect to the IP, $\sigma_{x_{IP}}$; (b) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
Figure 8: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of the Gaussian distribution of axial fluctuations of the luminometer position with respect to the IP, $\sigma_{z_{IP}}$; (b) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S-triangles, AS-diamonds), with the detector placed at the z-axis.
@@ 3.2 Beam related uncertainties
Several systematic uncertainties may arise from uncertainties of the properties of the beams and their delivery to the interaction point. The list of CEPC beam parameters addressed in this study is given in Table 1 Gao .
Table 1: CEPC TDR beam parameters at the $Z^{0}$ pole.
@@
Here we consider, again similarly to JINST :
@@
    maximal permanent bias ($\Delta E$) of a single beam energy with respect to the other beam, resulting in a longitudinal boost of the incoming $e^{+}e^{-}$ system with respect to the laboratory frame, and consequently, boost of the Bhabha center-of-mass system,
@@
    maximal RMS of the Gaussian distribution of the beam energy spread ($\sigma_{E_{BS}}$), responsible for longitudinal boost on event-by-event basis, leading to the overall count loss of order of $\mathrm{10^{-4}}$,
@@
    maximal radial ($\Delta x_{IP}^{BS}$) and axial ($\Delta z_{IP}^{SY}$) IP position displacements with respect to the luminometer arms, caused by the finite beam sizes (former) and beam synchronization (latter),
@@
    maximal time shift in beam synchronization ($\Delta\tau_{SY}$), determined from $\Delta z_{IP}^{SY}$, which may cause the IP longitudinal displacement $\Delta z_{IP}^{SY}$.
In the same manner as in the previous chapter, results for effects 1-3 are illustrated on Figures 9-12 and the upper limits, together with the upper limits of mechanical parameters from Section 3.1, are summarized in Table 2.
Table 2: Minimal absolute precision of luminometer mechanical parameters and beam parameters, each contributing as $10^{-4}$ to the relative uncertainty of $\mathcal{L}$ at the $Z^{0}$ pole. The average net center-of-mass energy uncertainty $\Delta E_{CM}$ limit is derived by error propagation from the Bhabha cross-section dependence on the center-of-mass energy.
@@
As can be seen from Table 2, mechanical precision of the luminometer’s inner aperture mechanical precision seems to be the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$. The longitudinal boost of the Bhabha center-of-mass frame with respect to the laboratory frame, $\beta_{z}$, may be caused by any difference in beam energies, $\delta E$ ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), in the form of bias, $\Delta E$, of one beam with respect to the other, or as a random asymmetry caused by the beam energy spread in the presence of other radiative losses (ISR and FSR). Limits of $\Delta E$ and $\sigma_{E_{BS}}$ given in Table 2 are derived from Figures 9(b) and 10(b) and rounded with respect to statistical size of the sample. From Table 2 it can be seen that in terms of the beam-related parameters precision, displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision significantly, while some dependencies between observables will be modified due to loss of spacial symmetries of Bhabha events.
@@
Figure 10: (b) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal RMS of the BES Gaussian distribution, corresponding to the relative systematic uncertainty of $\mathcal{L}$ of $10^{-4}$, $\sigma_{E_{BS}}$; (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S-triangles, AS-diamonds), with the luminometer on the z-axis.